# Let Me Think! A Long Chain-of-Thought Can Be Worth Exponentially Many Short Ones

### v1

@@ Results
After training models for the Bridge(5) task with Walk-L CoT strategies, we find that the accuracy of the models consistently increases with $L$, which shows that the models trained on more exploratory and longer walks perform better (See Figure 14).
@@
We find that small models trained on DFS-BT CoTs solve the task consistently, while small models trained on DFS CoTs fail to solve the Bridge tasks of larger depths (See Figure 11), which can be explained by the smaller model’s more limited expressivity.
@@ C.4 LLM Experimental Details
We used vllm and 2 A100 GPUs ($80$ GB of memory each) for inference. The experiments to make each plot took less than four hours each. Debugging and hyperparameter tuning took under 120 GPU hours. We constructed $32$ random labelings of the bridge graph, and then, using prompts of the form Figure 12, create CoTs of 4096 tokens. Depending the model, we added the appropriate special tokens to make the input prompt from the user, and to make the model use thinking mode during the CoT. Each model recommended using temperature $0.6$ for thinking, which we did. We used a custom logit processor to make the model substitute the end thinking token and the eos token with the token for "wait", inspired by [40]. Then we truncate the CoT at intervals evenly spaced by tokens, and append the end of thinking token, and “Answer: Node [start node label] is in the same connected component as node ” before using the model to find the logits for the next token. The model is considered correct if the logit for the correct node is higher than the logit for the incorrect node introduced in the initial prompt [^5] With some tie breaking when the logits are within $1\mathrm{e}{-8}$ of each other. We found that techniques weighting the confidence by the magnitude of the logits or their difference did not significantly change any results.. For parallel scaling, we generated up to $64$ distinct CoTs for each graph, and analytically calculated the probability that a random subsample would vote for the correct or incorrect solution (or tie). All of the results have a standard deviation of at most $0.08$.
@@ C.5 LLM Additional Experiments
In Figure 10 also tested the LLMs on a setting closer to the setting of Theorem 2, where the graph to be explored is two disjoint paths, and we once again confirm the theory, and see similar trends to those in Figure 1.
Figure 10: A comparison of parallel and sequential scaling for three LLMs tested on the $(s,t_{1},t_{2})$-connectivity problem for a graph that is the disjoint union of two paths. Note the similar trend to Figure 1.
@@
Figure 13: Number of samples to get larger than 95% best-of-N accuracy by parallel scaling models trained with different CoT strategies with temperature 1.0. The average over 5 runs is reported.

### v2

@@ Results
After training models for the Bridge(5) task with Walk-L CoT strategies, we find that the accuracy of the models consistently increases with $L$, which shows that the models trained on more exploratory and longer walks perform better (See Figure˜14).
@@
We find that small models trained on DFS-BT CoTs solve the task consistently, while small models trained on DFS CoTs fail to solve the Bridge tasks of larger depths (See Figure˜13), which can be explained by the smaller model’s more limited expressivity.
@@ C.4 LLM experimental details
For the AIME2024 experiment, we used H200 GPUs. Each run took approximately 1.5 hours, for a total of about 24 H200 GPU-hours. For the graph connectivity experiment, we used vllm and 2 A100 GPUs ($80$ GB of memory each) for inference. The experiments to make each plot took less than four hours each. Debugging and hyperparameter tuning took under 120 GPU hours. We constructed $32$ random labelings of the bridge graph, and then, using prompts of the form Figure˜11, create CoTs of 4096 tokens. Depending the model, we added the appropriate special tokens to make the input prompt from the user, and to make the model use thinking mode during the CoT. Each model recommended using temperature $0.6$ for thinking, which we did. We used a custom logit processor to make the model substitute the end thinking token and the eos token with the token for "wait", inspired by [40]. Then we truncate the CoT at intervals evenly spaced by tokens, and append the end of thinking token, and “Answer: Node [start node label] is in the same connected component as node ” before using the model to find the logits for the next token. The model is considered correct if the logit for the correct node is higher than the logit for the incorrect node introduced in the initial prompt [^4] With some tie breaking when the logits are within $1\mathrm{e}{-8}$ of each other. We found that techniques weighting the confidence by the magnitude of the logits or their difference did not significantly change any results.. For parallel scaling, we generated up to $64$ distinct CoTs for each graph, and analytically calculated the probability that a random subsample would vote for the correct or incorrect solution (or tie). All of the results have a standard deviation of at most $0.08$.
@@ C.5 LLM additional experiments
In Figure˜10 also tested the LLMs on a setting closer to the setting of Theorem˜2, where the graph to be explored is two disjoint paths, and we once again confirm the theory, and see similar trends to those in Figure˜1.
Figure 10: A comparison of parallel and sequential scaling for three LLMs tested on the $(s,t_{1},t_{2})$-connectivity problem for a graph that is the disjoint union of two paths. Note the similar trend to Figure˜1.
@@ Appendix D Further related work
Figure 12: A comparison of parallel and sequential scaling for s1-32B [40] on AIME2024 [44]. For parallel scaling, answers are sampled with temperature 1.0 and aggregated by majority vote.